OpenStudy (anonymous):

Which medium (the glass, water or diamond), is the most dense and how do you know?

OpenStudy (anonymous):

Diamond

OpenStudy (s3a):

For something to be more dense, it has to have more mass per some equivalent volume. So, if you could get everything you want to compare to be in the same shape and size then, you could figure out its mass by weighing it and whichever medium/object has the largest mass (and the same volume) has the largest density.

OpenStudy (s3a):

When you figure out the weight, you can figure out the mass by using F = -mg where F is the weight (since weight is a force) and g is the gravitational acceleration which is assumed to be constant for simplicity since it varies very slightly around different parts of Earth. If you're class hasn't reached this level of complexity yet, all you need to take away is my previous post and the fact that knowing the weight of something allows you to easily determine its mass. Mass is a measure of how much matter makes up a certain object whereas weight is how much gravitational force acts on that matter.

OpenStudy (shane_b):

In addition to what @s3a said: \[density=\frac{mass}{volume}\]You can measure the mass and volume of each object and determine the density as shown in the formula above: For the materials listed in the problem: \[Diamond = \frac{3.5g}{cm^3}\]\[Glass = \frac{2.5g}{cm^3}\]\[Water = \frac{1g}{cm^3}\]Note that the density of each can vary somewhat depending on several factors (type of glass, temperature, etc) but the above values are close enough for a comparison...assuming sample is homogenous.

OpenStudy (s3a):

Oops, I forgot to mention density = mass/volume when I wrote my messages! It's good you did, @Shane_B!

OpenStudy (anonymous):

optical density or physical density?

OpenStudy (anonymous):

if optical density the one with the greatest refractive index diamond